OpenStudy (zootedyoungin):

My equation is There are 6 apples on the floor, X amount of dogs, the amount of dogs and apples together is 12, how many dogs are there?

Mehek (mehek14):

6 apples x dogs total is 12 6 + x = 12 All you have to do is isolate x, so subtract by 6 on each side.

OpenStudy (zootedyoungin):

Hmm

OpenStudy (zootedyoungin):

6 - 12 = -6?

OpenStudy (zootedyoungin):

subtract 6 on both sides

OpenStudy (zootedyoungin):

is anyone here good with banking

OpenStudy (partyshotz2409):

12-6=x

Mehek (mehek14):

What Party said, 12 - 6 = answer

OpenStudy (zootedyoungin):

lol...

OpenStudy (partyshotz2409):

can someone check out my question

OpenStudy (zootedyoungin):

x = 6

OpenStudy (zootedyoungin):

how do i gain smart score

OpenStudy (partyshotz2409):

good

OpenStudy (partyshotz2409):

you answer questions and talk with people on your questions

OpenStudy (zootedyoungin):

i need 2 smart score ._.

OpenStudy (zootedyoungin):